• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 01:29:01 on August13,2006

    Entry number 179517

    keyword=status

    Here is the grim status:

    1. L-HRS dipole helium level remains 1.2%. Can't turn on the magnet
    until it is 80%. We are stuck.

    2. Target at empty. We reduce JT to reduce cryo load to 9 g/sec.

    3. 1 uA in hall to keep the accelerator "warm".

    4. There is a water puddle on Hall A floor in the vicinity of Q2.
    It does not appear to be growing. I guess its related to the cryo
    problems earlier.